# GenOM: Ontology Matching with Description Generation and Large Language Model

### v1

@@ 1 Introduction
In recent years, the rapid growth of domain-specific ontologies has led to a growing need for semantic interoperability and knowledge integration across diverse knowledge systems [1]. Ontologies are developed to formally represent concepts and relationships in a domain, yet they are often constructed in isolation, following distinct modelling choices, terminological conventions, and structural assumptions. This independence has resulted in significant heterogeneity across ontologies, which poses considerable challenges to the integration and reuse of knowledge.
Ontology Matching (OM) as known as ontology alignment, the task of identifying semantic correspondences between entities in different ontologies, has therefore become a crucial area of research [2, 3]. By establishing links such as equivalence (indicating that two concepts represent the same or highly similar meaning) or subsumption (where one concept is a more general or specific variant of the other), ontology alignment facilitates accurate knowledge translation and consistent information exchange between systems. However, matching concepts across ontologies is far from straightforward. Three major sources of heterogeneity commonly hinder this process: (1) Terminological differences, where the same concept may be described using different labels or synonyms; (2) Structural differences, reflecting the varying levels of complexity in ontology design—from deeply nested hierarchies to flat enumerative lists; (3) Granularity differences, where the same domain knowledge may be captured with differing levels of detail or abstraction.
These variations significantly increase the cognitive and computational burden of OM. The challenge is further magnified by the rapidly growing scale of modern ontologies. For example, SNOMED-CT [4], a widely adopted clinical terminology, contains several hundred thousand medical concepts. As ontologies continue to expand in size and complexity, manual alignment methods become increasingly infeasible, underscoring the necessity for automated or semi-automated alignment techniques capable of operating at scale.
Traditional OM systems, such as LogMap [5] and AML [6], primarily rely on string matching, indexing and structure matching technologies. They often fall short in capturing or fully utilising the underlying semantic information of concepts. With the advent of large language models (LLMs), their remarkable capabilities in text understanding and generalisation have attracted significant attention. Recently, several ontology alignment systems have begun to incorporate LLMs to better capture the semantics of complex concepts across heterogeneous ontologies. These approaches typically leverage LLMs for identifying semantic similarities between concept pairs via embedding, and/or making direct alignment decisions via generation. For example, LLM4OM [7] employs ChatGPT and OpenAI embeddings for pairwise matching (see Section 2 for more related work analysis). However, current LLM-based OM approaches still exhibit notable limitations. Some of them, despite leveraging powerful language models, struggles to deliver satisfactory performance on more complex OM tasks. Some others can achieve promising results on certain benchmarks, but they may rely on LLMs with very large-scale parameters (e.g., 70B LLaMA-2 used in Olala [8]) that imposes substantial computational demands, raising concerns about scalability.
To address these limitations, we propose a novel ontology matching framework named GenOM which utilises LLMs and extended textual definitions of concepts. The framework begins by extracting both lexical and structural information from the source and target ontologies. This information is then semantically enhanced using an LLM, resulting in more informative and context-aware concept descriptions. Subsequently, these enriched representations are embedded into vector space, enabling the retrieval of candidate alignments based on semantic similarity. A lightweight 7B-parameter LLM is employed to assess the equivalence of candidate pairs through a classification-based approach, while traditional exact matching techniques are incorporated to supplement and refine the alignment results. The effectiveness of GenOM is demonstrated through comprehensive experiments evaluated using standard metrics such as precision, recall, F1-score, mean reciprocal rank (MRR), and Hit@K. The model achieves competitive performance compared to several state-of-the-art ontology alignment systems, highlighting its robustness and practical applicability. Extensive ablation studies were also conducted to demonstrate the effectiveness of the proposed framework across multiple dimensions.
@@ 2 Related Work
At present, OM approaches can be broadly categorised into four main types: traditional knowledge-based systems, machine learning-based systems, pre-trained language model-based systems and more recently, large language models (LLMs)-based systems. Traditional systems, such as LogMap and AML, rely primarily on lexical similarity, structural heuristics, and external resources like UMLS or WordNet [9]. LogMap [5] extracts class names and searches for matches via external lexicons while addressing logical inconsistencies by selecting alignments with higher confidence scores. AML [6] employs multiple matching strategies, including exact and character-based matchers, and has shown strong performance on medical datasets. However, these systems depend heavily on curated lexicons and often fail to fully capture and utilise complex semantics of different kinds.
@@
With the rise of machine learning, several approaches have emerged that model alignment as a classification or (embedding-based) similarity learning task [10, 11, 12, 13, 14, 15]. For instance, DeepAlignment [11] vectorises class names and computes Euclidean distances to assess similarity, while the CNN-based system [15] use character-level embeddings and hierarchical context to train binary classifiers for equivalence detection. Although these methods offer improvements over traditional techniques, they often require large annotated datasets and extensive parameter tuning. Moreover, their domain-specific nature limits transferability across ontologies in different fields.
@@
Using encoder-based pre-trained language models like BERT  [16] have leveraged the representational power of contextual embeddings and a memorization based on large-scale parameters learned from corpora to address some of these limitations. BioSTransformers [17] adopt a Siamese architecture based on domain-specific BERT models to compute semantic similarity between biomedical concepts. Built on the BERT architecture [18], BERTMap fine-tunes a domain-specific BERT model on ontology alignment corpora, enabling it to capture subtle semantic differences even when lexical overlap is low. This approach has demonstrated strong results, particularly in biomedical applications. BERTSubs [19] takes a similar architecture as BERTMap but focuses on the subsumption relationship.
More recently, LLMs such as GPT-3.5, GPT-4, and T5-XXXL have been applied to ontology alignment, offering stronger generalisation and semantic reasoning capabilities. Several studies have explored prompt-based querying and retrieval-augmented generation to support alignment tasks. For example, Norouz et al. [20] used GPT-4 to align ontology via prompts, observing high recall but reduced precision due to the model incorrectly classifying subclass relationships as equivalence relations. Yuan et al. [21] tested both open- and closed-source models on medical alignment tasks, using structural context to enhance predictions. Other work has introduced hybrid frameworks combining vector similarity retrieval (e.g., using SBERT) with LLM verification stages [7, 22], aiming to reduce hallucination and improve alignment quality. The Olala system [8] further integrates embedding-based candidate filtering and post-processing with LLaMA-2 for final alignment decisions. While LLM-based approaches have shown considerable potential, they still face challenges including scalability to large ontologies such as SNOMED CT, struggles to deliver satisfactory performance on more complex OM tasks and computational cost, particularly when relying on closed-source or very large models.
@@ 3.1 Task Formulation
The OM task can be formally defined as follows. Given two ontologies, referred to as the source ontology $O_{\text{s}}$ and the target ontology $O_{\text{t}}$, let $C_{s}$ denote the set of named concepts in $O_{\text{s}}$ and $C_{t}$ denote the set of named concepts in $O_{\text{t}}$. The objective is to identify a set of mappings, where each mapping consists of a concept pair $(c_{s},c_{t})$, with $c_{s}\in C_{s}$ and $c_{t}\in C_{t}$, that are considered semantically related. Formally, the alignment output is represented as:
$M=\left\{(c_{s},c_{t},\alpha)\mid c_{s}\in C_{s},\,c_{t}\in C_{t},\,\alpha\in[0,1]\right\}$ (1)
where $\alpha$ denotes the confidence score, quantifying the degree of semantic equivalence between $c_{s}$ and $c_{t}$.
This score often serves as a basis for selecting high-confidence mappings.
3.2 System Architecture
As shown in Figure 1, GenOM comprises five main components:
  1. 1.
    Ontology Data Extraction: Structural and lexical information is extracted for each named concept from both the source and target ontologies. This includes labels, synonyms, parent concepts, and axioms of concept equivalence.
  2. 2.
    Definition Generation: An LLM is prompted to generate natural language definitions or paraphrased descriptions of each concept, based on the extracted information. This step enhances the semantic representation of concepts, especially for those lacking explicit textual definitions.
  3. 3.
    Candidate Mapping Generation: Using the embedding model, pairwise cosine similarity scores are computed between the source and target concepts to generate top-$k$ candidate mappings.
  4. 4.
    LLM-Based Equivalence Judgement: For each candidate mapping, an LLM is queried to determine whether the mapping’s two concepts are semantically equivalent, with prompts that incorporate enriched definitions and structural context.
  5. 5.
    Post-processing and Result Fusion: Filtering is performed based on both the probability distribution of the LLM outputs and the cosine similarity scores, retaining only high-confidence alignment results. In addition, exact matching modules are applied to recall highly confident matches based on identical labels.
By concept definition generation, candidate retrieval, and alignment judgement, GenOM integrates the strengths of embedding-based similarity, exact lexical matching, and LLM reasoning.
@@ 3.2 System Architecture
3.2.1 Ontology Data Extraction
Both lexical and structural characteristics of each concept from the source and target ontologies are extracted and utilized to support the following alignment process. Specifically, the extracted information includes the concept’s label (defined by the annotation property rdfs:label), a set of synonyms (retrieved using the annotation properties listed in Table 1), and its parent concepts. For concepts defined using EquivalentClass axioms, the built-in verbalisation module in DeepOnto [23] is used to convert logical expressions into natural language descriptions (Table 2).
@@ 3.2.1 Ontology Data Extraction
Description Logic Axiom
@@
Verbalized Description
@@
Table 2: Example of DL Axiom and Its Verbalized Description
3.2.2 Definition Generation
LLMs encode extensive general purpose and domain knowledge. They are leveraged and integrated with concept information explicitly represented in the ontology to enhance the semantic representation of the concept. In this framework, definitions for medical concepts are generated by prompting an LLM with background information extracted from the ontology, including the concept’s label, synonyms, parent concepts, and, where applicable, the verbalised descriptions of logical expressions from EquivalentClass axioms.
Table 3 presents the prompt template employed for definition generation. In this template, the LLM is provided with both the available concept-specific information and the name of the source ontology. This additional context helps the LLM recall relevant domain knowledge encoded in its parameters, thereby generating definitions that are more accurate and context-aware. The inclusion of the ontology name in the prompt acts as supplementary guidance, particularly for sparsely annotated concepts.
@@ 3.2.2 Definition Generation
Output:
  A medicinal product specifically formulated to contain solely betamethasone and calcipotriol as its active ingredients, designed for the treatment or management of specific dermatological conditions.
@@
3.2.3 Candidate Mapping Generation
@@
To generate candidate concept pairs for alignment, this framework adopts an embedding-based retrieval strategy. Concepts from both the source ontology $O_{\text{s}}$ and the target ontology $O_{\text{t}}$ are first encoded into fixed-size vector representations. Each concept is represented using a combination of its label, synonyms, and its enriched definition produced in the previous step. Figure 2 shows an example of the input text fed into the embedding model. Structural information such as hierarchical relations is deliberately excluded at this stage to reduce complexity in embedding. The framework adopts the text-embedding-3-small [^1] https://platform.openai.com/docs/models/text-embedding-3-small model for embedding generation. Compared to traditional encoder-based models such as Sentence-BERT [24], this LLM-based embedding model demonstrates superior capability in distinguishing complex biomedical concepts.
Once the embeddings are obtained, a cosine similarity-based retrieval process is applied to identify, for each source concept, the top-$k$ most semantically similar concepts from the target ontology. Candidate selection is based on vector similarity, allowing the system to retrieve a shortlist of potentially equivalent concept pairs. These candidates are subsequently passed to the next stage for semantic equivalence assessment.
@@
Embedding Input Example Label: Product containing only betamethasone and calcipotriol (medicinal product); Synonyms: Betamethasone and calcipotriol only product; Definition: A medicinal product specifically formulated to contain solely betamethasone and calcipotriol as its active … ;
Figure 2: Concept information input format for embedding.
3.2.4 LLM-Based Equivalence Judgement
@@
In this stage, an LLM is employed to determine whether each candidate concept pair represents a semantic equivalence. Rather than prompting the model to generate full descriptive justifications, which would be time-consuming and potentially verbose, a lightweight classification strategy is adopted. Specifically, each concept pair is presented via a prompt designed to elicit a binary response — YES if the concepts are equivalent, and NO otherwise.
To support this, a prompt (as shown in Table 4) is constructed with strong instructional guidance, encouraging the model to respond using only a single classification token.
The predicted equivalence score is then computed based on the probability of the YES token, extracted directly from the model’s output logits. Specifically, given the model’s output logits $\mathbf{z}\in\mathbb{R}^{V}$ at the final decoding position (where $V$ is the vocabulary size), the softmax function is applied to convert the logits into a probability distribution:
$P(\texttt{YES})=\text{softmax}(\mathbf{z})_{\texttt{YES}}=\frac{\exp(z_{\texttt{YES}})}{\sum_{i=1}^{V}\exp(z_{i})}$
@@
Here, $z_{\texttt{YES}}$ denotes the logit corresponding to the token YES. The resulting probability serves as the model’s confidence in semantic equivalence for a given concept pair. Concept pairs with $P(\texttt{YES})$ exceeding a predefined threshold are retained for alignment.
This probability-based scoring approach significantly reduces inference time and simplifies decision-making, as it avoids generating full-length text responses and instead relies on a single-token classification strategy, while maintaining high alignment precision.
@@
Table 4: The prompt template used for binary equivalence classification between ontology concepts. Dynamic fields are populated with structured ontology information.
3.2.5 Post-processing and Result Fusion
To ensure the quality and reliability of the final alignment output, a post-processing stage is applied to filter and refine the results generated by the LLM. First, a threshold $\lambda_{\text{prob}}$ is imposed on the token-level probability associated with the YES response. Candidate pairs with confidence scores below this threshold are discarded. In parallel, the cosine similarity scores obtained during candidate generation are also considered, and pairs with lower than $\lambda_{\text{cs}}$ embedding similarity are removed to prevent semantically distant matches from being retained.
After this dual-filtering step, To further enhance precision, outputs from the LLM module were merged with the results of two exact matching systems, LogMapLt [^2] https://github.com/ernestojimenezruiz/logmap-matcher and BERTMapLt [^3] https://github.com/KRR-Oxford/DeepOnto/tree/main/src/deeponto/align/bertmap, which are lightweight versions of their original models, both simplified to include only the string matching component. This fusion combines semantic reasoning with surface-level matching, improving overall coverage while preserving precision. The resulting set constitutes the final alignment output.
4 Experiment and Evaluation

### v2

@@ 1 Introduction
In recent years, the rapid growth of domain-specific ontologies has intensified the demand for semantic interoperability and knowledge integration across heterogeneous information systems [1]. In particular, the biomedical domain [2] has witnessed the proliferation of large and complex ontologies, such as those used to model diseases, anatomy, and clinical procedures, which further amplifies this challenge. While ontologies serve as structured and machine-interpretable representations of conceptual knowledge, they are often engineered independently by different research communities or organisations. Each ontology typically reflects particular domain-specific objectives, modelling assumptions, and terminological conventions, which collectively result in substantial structural and semantic heterogeneity. This diversity, though valuable for capturing local perspectives, poses significant challenges to semantic interoperability and the effective integration of knowledge across heterogeneous information systems. Consequently, bridging conceptual discrepancies among ontologies has become a central challenge for achieving seamless knowledge exchange and reuse in data-intensive environments. Against this backdrop, the goal of this paper is to develop an ontology matching (OM) framework that can effectively bridge heterogeneous ontologies by leveraging the reasoning and generation capabilities of large language models (LLMs).
As a core mechanism for achieving this goal, ontology matching (OM) has emerged as a pivotal research area in response to the growing need for semantic integration across heterogeneous knowledge bases [3, 4]. Its primary objective is to uncover semantic correspondences between entities that originate from different ontologies, thereby enabling the consistent interpretation and exchange of knowledge. Through establishing formal relations such as equivalence, where two entities express the same or highly similar meanings, and subsumption, where one entity denotes a more general or more specific concept, OM serves as a cornerstone for semantic interoperability in distributed information environments.
Nevertheless, achieving accurate alignments remains highly challenging due to the intrinsic heterogeneity of ontologies. This heterogeneity is manifested in several forms. First, terminological variation arises when identical concepts are described using distinct lexical labels or synonyms. Second, structural variation reflects the diversity of ontology design choices, ranging from deep hierarchical taxonomies to flat enumerative structures. Third, granularity variation occurs when similar knowledge is modelled at differing levels of abstraction. Together, these inconsistencies amplify both the cognitive and computational complexity of the matching task. The challenge becomes even more pronounced as ontologies continue to scale. For example, SNOMED CT [5], one of the largest biomedical terminologies, comprises hundreds of thousands of concepts, which has led some studies to further decompose it into smaller, task-specific subontologies for improved tractability [6, 7], yet purely manual alignment remains impractical at this scale, further motivating the need for scalable automated solutions.
Over the past decade, a range of automated systems, such as LogMap [8] and AgreementMakerLight (AML) [9], have been proposed to address this challenge. These approaches integrate lexical similarity, structural heuristics, and background knowledge resources to identify candidate mappings. While effective in well-structured and lexically rich scenarios, they often fail to capture the deeper semantics encoded in complex concept descriptions. Recent progress in LLMs has fundamentally reshaped this landscape. By virtue of their powerful contextual reasoning and linguistic generalisation abilities, LLMs offer an unprecedented opportunity to represent and compare ontological knowledge at a semantic level. Several contemporary OM systems have begun to incorporate LLMs at different stages of the alignment pipeline, including semantic embedding and direct alignment prediction. For instance, LLM4OM [10] integrates ChatGPT with OpenAI embeddings to support pairwise concept matching, while Olala [11] leverages LLM-based semantic reasoning to assess concept equivalence. A more detailed review is given in Section 2. Despite these advances, current LLM-based approaches remain constrained by practical and methodological limitations. Some rely on extremely large models (e.g., the 70B-parameter LLaMA-2 employed in Olala [11]), which incur considerable computational costs and hinder scalability, whereas others [12] struggle to maintain robustness when confronted with complex or large-scale alignment scenarios.
To address these limitations, we develop GenOM, an ontology matching framework that leverages LLMs for semantic enrichment, candidate generation, and mapping judgement. [^1] A preliminary version of this idea, focusing on LLM-based definition generation, was presented at the OM Workshop @ ISWC 2025 [13]. GenOM begins by extracting lexical and structural information from both source and target ontologies—including labels, synonyms, parent relations, and equivalence axioms—and employs an LLM-driven description generation module to produce semantically enriched textual definitions for each concept. These enriched descriptions are then embedded into a vector space to support efficient retrieval of semantically similar candidates, while an LLM-based mapping judgement component and a lightweight exact-matching module jointly refine the alignments.
In this work, we instantiate GenOM using a diverse suite of LLMs, including Qwen2.5-7B-Instruct, Qwen2.5-14B-Instruct, Qwen2.5-32B-Instruct, and Llama3-8B-Instruct, which vary in architectural design and parameter scale. This enables a systematic analysis of how model capacity influences definition quality, semantic representation, and overall alignment performance.
We conduct a comprehensive experimental study across five OM tasks involving biomedical ontologies of SNOMED-CT, NCIT, FMA, ORDO, DOID, and OMIM (see Section 4.1 for more details). Our evaluation examines GenOM from four perspectives: (1) its overall alignment performance is compared with strong existing systems; (2) the quality of LLM-generated definitions is assessed using lexical metrics and an LLM-as-a-judge semantic evaluation; (3) performance differences across LLMs of varying scales and architectures are analysed; (4) ablation studies are conducted to isolate the contributions of semantic enrichment, candidate generation, and mapping judgement. These experiments collectively demonstrate the robustness of GenOM and the importance of LLM-driven semantic enrichment in ontology matching. An open-source implementation of the proposed approach is available. [^2] Code repository: https://github.com/AlanS87/genom-pipeline
Contributions. This work makes the following key contributions:
  • GenOM: an LLM-enhanced ontology matching framework. We introduce GenOM, a new ontology matching framework that integrates LLM-driven semantic enrichment with embedding-based candidate generation and LLM-based mapping judgement. GenOM achieves strong performance across multiple biomedical alignment tasks, demonstrating the effectiveness of combining LLM semantics with traditional matching components.
  • A novel evaluation methodology for assessing the quality of LLM-generated concept definitions. We propose a dedicated evaluation scheme for concept definitions in OM tasks, combining lexical metrics, alignment-based measures, and an LLM-as-a-judge semantic assessment to quantify the correctness and discriminativeness of generated definitions.
  • A comparative analysis of LLMs of different scales for ontology matching. We systematically compare LLMs over varying parameter sizes and architectures, revealing how model capacity influences definition quality, threshold behaviour, and final alignment performance, and highlighting their potential for future ontology matching research.
The remainder of this paper is organised as follows. Section 2 reviews related work on ontology matching. Section 3 introduces the GenOM framework in detail, including its overall methodology and the design of each core component. We report the performance of GenOM on multiple biomedical ontology matching tasks, assess it from multiple perspectives, and compare it against several strong baseline systems in Section 4. Section 5 presents an in-depth analysis of the quality of LLM-generated definitions, examining their correctness and discriminative capacity. Section 6 investigates the impact of different LLMs, analysing how model scale and architecture influence alignment behaviour and performance. Section 7 provides an ablation study to analyse the impact of LLM-generated definitions on different stages of the alignment process. Finally, Section 8 concludes the paper with a discussion of the main findings, limitations, and directions for future research.
@@ 2 Related Work
At present, OM approaches can be broadly categorised into four main types: traditional knowledge-based systems, machine learning-based systems, pre-trained language model-based systems and more recently, large language models (LLMs)-based systems. Traditional systems, such as LogMap and AML, rely primarily on lexical similarity, structural heuristics, and external resources like UMLS or WordNet [14]. LogMap [8] extracts class names and searches for matches via external lexicons while addressing logical inconsistencies by selecting alignments with higher confidence scores. AML [9] employs multiple matching strategies, including exact and character-based matchers, and has shown strong performance on medical datasets. However, traditional ontology matching systems largely rely on string-based similarity measures and curated external lexicons, which have been shown to limit their coverage when semantically equivalent concepts differ substantially at the lexical level [15]. As a result, such approaches may fail to identify valid correspondences that are not explicitly recorded in external dictionaries or lexicons.
@@
With the rise of machine learning, several approaches have emerged that regard alignment as a classification or (embedding-based) similarity learning task [16, 17, 18, 19, 20, 21]. For example, DeepAlignment [17] vectorises class names and computes Euclidean distances to assess similarity, while the CNN-based system [21] uses character-level embeddings and hierarchical context to train binary classifiers for equivalence detection.
@@
In parallel, encoder-based pre-trained language models such as BERT [22] have been explored for ontology matching, leveraging contextualised representations learned from large-scale corpora. BioSTransformers [23] adopt a Siamese architecture based on domain-specific BERT models to compute semantic similarity between biomedical concepts. BERTMap [15] fine-tunes a domain-specific BERT model on ontology alignment corpora mainly automatically extracted from synonyms, enabling it to capture subtle semantic differences even when lexical overlap is low. This approach has demonstrated strong results, particularly in biomedical applications. BERTSubs [24] takes a similar architecture as BERTMap but focuses on the subsumption relationship.
More recently, the emergence of generative LLMs, including the GPT series, the Llama series, and the T5 series, has stimulated a new line of research in ontology alignment. These LLMs provide strong contextual reasoning and linguistic generalisation capabilities, making them attractive tools for capturing semantic relations that go beyond lexical similarity. A number of studies have explored prompting strategies and retrieval-augmented pipelines to exploit these capacities. For example, Norouz et al. [12] employed GPT-4o in a prompt-driven alignment workflow, reporting strong recall, while highlighting challenges in reliably distinguishing equivalence from hierarchical relations. Yuan et al. [25] evaluated both open- and closed-source LLMs on biomedical alignment tasks, incorporating structural cues to strengthen predictive reliability. Hybrid frameworks that combine embedding-based similarity retrieval with LLM verification have also been proposed [10, 26], aiming to mitigate hallucination and improve alignment robustness. The Olala system [11] further integrates embedding-based filtering and post-processing modules with Llama-2, which is responsible for generating final alignment decisions. Hu et al. [27] have explored enriching ontological information by leveraging LLMs to expand or clarify incomplete concept metadata. Their approach employs LLaMA-3.3-70B-Instruct to, for instance, transform abbreviated forms into their likely full expressions and to generate explanatory descriptions for concepts with limited textual context.
Traditional ontology matching systems have also begun to incorporate LLMs to enhance their existing pipelines. For instance, the latest version of LogMap [28] integrates GPT-4-mini and models from the Gemini family to assist in resolving uncertain and ambiguous candidate mappings. This hybrid integration enables the system to refine difficult alignment cases that conventional lexical- or structure-based heuristics struggle to disambiguate, resulting in modest improvements over the original LogMap pipeline.
Despite recent advances, existing LLM-based and LLM-augmented ontology matching approaches continue to face notable limitations. A central challenge arises from the fragmented and often sparse lexical information available within most ontologies [11, 10, 12]. LLMs must rely heavily on labels, short synonyms, and minimally structural cues, which provide insufficient semantic context for distinguishing between closely related or hierarchically similar medical concepts. As a result, current LLM-based systems often struggle in fine-grained disambiguation scenarios [10], where subtle terminological or relational differences must be accurately interpreted.
To address this limitation, our approach introduces an LLM-driven definition generation module that enriches each concept with a coherent, semantically informative description prior to alignment. These enriched definitions provide the LLM with a more holistic understanding of the underlying concepts, enabling more reliable and fine-grained mapping judgements compared with relying solely on the ontology’s native lexical fragments.
@@ 3.1 Task Formulation
OM is the process of discovering semantic correspondences between entities that belong to different ontologies. Formally, given two ontologies, the source ontology $O_{\text{s}}$ and the target ontology $O_{\text{t}}$, let $C_{\text{s}}$ and $C_{\text{t}}$ denote sets of named concepts respectively contained in $O_{\text{s}}$ and $O_{\text{t}}$. The goal of OM is to identify a set of mappings that link semantically related concepts across the two ontologies. Each mapping consists of a concept pair $(c_{\text{s}},c_{\text{t}})$, where $c_{\text{s}}\in C_{\text{s}}$ and $c_{\text{t}}\in C_{\text{t}}$, together with an associated confidence value $\alpha_{st}\in[0,1]$, which quantifies the strength or likelihood of their semantic equivalence. The resulting alignment can therefore be expressed as a set of scored correspondences:
$M=\left\{(c_{\text{s}},c_{\text{t}},\alpha_{st})\mid c_{\text{s}}\in C_{\text{s}},\,c_{\text{t}}\in C_{\text{t}},\,\alpha_{st}\in[0,1]\right\}$ (1)
A higher value of $\alpha_{st}$ indicates greater semantic proximity between $c_{\text{s}}$ and $c_{\text{t}}$, and is often used as a confidence score for ranking candidate mappings or filtering high-confidence alignments. In practical settings, the OM task can encompass various types of semantic relations, such as equivalence, subsumption, and relatedness, depending on the intended application or evaluation protocol. Accordingly, the OM task can be regarded as a process of semantic reconciliation that transforms heterogeneous conceptual representations into a shared and interoperable knowledge space.
In this study, the focus is restricted to the equivalence problem, where the objective is to determine concept pairs that denote an identical meaning across the source and target ontologies.
3.2 The Workflow of the GenOM Framework
The overall workflow of GenOM, illustrated in Figure 1, comprises five main components. 1) Ontology data extraction collects structural and lexical information for each named concept in the source and target ontologies, including labels, synonyms, parent concepts, and equivalence axioms. 2) An LLM-based definition generation module produces natural language definitions based on paraphrased descriptions of concepts using the extracted information. This enriches their semantic representations—particularly for entities without explicit textual descriptions [29, 30]. 3) A candidate generation stage employs an embedding-based retrieval process to compute pairwise cosine similarities between source and target concepts, thereby identifying the top-$k$ candidate mappings. 4) An LLM-based equivalence judgement module evaluates each candidate pair, determining whether the two concepts convey equivalent meanings by leveraging their generated definitions and structural context. 5) A post-processing and result fusion stage integrates the LLM prediction scores with cosine similarities and exact lexical matching results, retaining only high-confidence correspondences.
Through the integration of definition generation, similarity retrieval, and semantic reasoning, GenOM combines the complementary strengths of embedding-based similarity, lexical precision, and LLM-based contextual understanding to achieve robust and scalable ontology matching.
@@ 3.2 The Workflow of the GenOM Framework
1)Ontology Data Extraction
Both lexical and structural characteristics of each concept from the source and target ontologies are extracted and utilized to support the following alignment process. Specifically, the extracted information includes the concept’s label (defined by the annotation property rdfs:label), a set of synonyms (retrieved using the annotation properties listed in Table 1), and its immediate parent concepts as defined by the rdfs:subClassOf relation.
For concepts defined using EquivalentClass axioms, the built-in verbalisation module in DeepOnto [31] is used to convert logical expressions into natural language descriptions (Table 2).
@@
EquivalentClass Axiom in Description Logic
@@
Verbalized Description by DeepOnto Toolkit
@@
Table 2: Example of EquivalentClass axiom, and the verbalized description of the logical expression
2)Definition Generation
LLMs encode extensive general purpose and domain knowledge. They are leveraged and integrated with concept information explicitly represented in the ontology to enhance the semantic representation of the concept. In GenOM, definitions for medical concepts are generated by prompting an LLM with background information extracted from the ontology, including the concept’s label, synonyms, parent concepts, and, where applicable, the verbalised descriptions of logical expressions from EquivalentClass axioms.
Table 3 presents the prompt template employed for definition generation. In this template, the LLM is provided with both the available concept-specific information and the name of the source ontology. This additional context is intended to provide the LLM with high-level domain cues, which may facilitate the generation of more contextually appropriate definitions. The ontology name serves as supplementary guidance rather than an explicit source of supervision, and is particularly useful when concept-level annotations are sparse.
@@
3)Candidate Mapping Generation
@@
To generate candidate concept pairs for alignment, GenOM adopts an embedding-based retrieval strategy. Concepts from both the source ontology $O_{\text{s}}$ and the target ontology $O_{\text{t}}$ are first encoded into fixed-size vector representations. Each concept is represented using a combination of its label, synonyms, and its enriched definition produced in the previous step. Figure 2 shows an example of the input text fed into the embedding model. Structural information such as hierarchical relations is deliberately excluded at this stage to reduce complexity in embedding. The framework adopts the text-embedding-3-small [^3] https://platform.openai.com/docs/models/text-embedding-3-small model provided by OpenAI for embedding generation. Compared to traditional encoder-based models such as Sentence-BERT [32], this embedding model benefits from large-scale pre-training on diverse textual corpora and is better suited to capturing nuanced semantic differences expressed in longer, definition-enriched inputs, which is particularly important for distinguishing fine-grained biomedical concepts.
After generating embeddings for all target ontology concepts, we build a Hierarchical Navigable Small World (HNSW) index using the FAISS library [33]. The index is constructed in the cosine similarity space by normalising all vectors to unit length, allowing FAISS to treat cosine similarity as equivalent to inner-product distance. This HNSW structure enables efficient approximate nearest-neighbour retrieval with sublinear query complexity, substantially reducing the computational cost compared with exhaustive pairwise similarity computation.
Once the index is constructed, each source concept embedding is queried against the FAISS HNSW index to retrieve its top-$k$ most similar target concepts under cosine similarity. This indexing-based retrieval mechanism significantly narrows the search space and provides a compact set of candidate mappings, which are subsequently passed to the LLM-based mapping judgement component.
Concept Representation Template Label: {label_string} Synonyms: {synonym_1}; {synonym_2}; … ; {synonym_n} Definition: {enriched_definition_generated_by_LLM}
@@
Embedding Input Example Label: Product containing only betamethasone and calcipotriol (medicinal product) Synonyms: Betamethasone and calcipotriol only product Definition: A medicinal product specifically formulated to contain solely betamethasone and calcipotriol as its active …
Figure 2: Template and example of concept information used for embedding-based representation.
4)LLM-based Equivalence Judgement
@@
In this stage, an LLM is employed to determine whether each candidate concept pair represents a semantic equivalence. Rather than prompting the model to generate full descriptive justifications, which would be time-consuming and potentially verbose, a lightweight classification strategy is adopted. Specifically, each concept pair is presented via a prompt designed to elicit a binary response — YES if the concepts are equivalent, and NO otherwise. To support this, a prompt (as shown in Table 4) is constructed with strong instructional guidance, encouraging the model to respond using only a single classification token.
The predicted equivalence score is computed using the probability assigned to the YES token, extracted directly from the model’s output logits. This follows the probabilistic interpretation of neural network outputs widely adopted in modern confidence estimation research, where token-level softmax probabilities are treated as calibrated or approximately calibrated confidence indicators [34, 35, 36].
Concretely, since the mapping judgement is explicitly formulated as a binary decision, we restrict the probability computation to the set of valid output labels $\mathcal{Y}=\{\texttt{YES},\texttt{NO}\}$. Given the model’s output logits $z_{\texttt{YES}}$ and $z_{\texttt{NO}}$ at the final decoding position, the probability of semantic equivalence is computed as:
$P(\texttt{YES})=\text{softmax}(\mathbf{z})_{\texttt{YES}}=\frac{\exp(z_{\texttt{YES}})}{\exp(z_{\texttt{YES}})+\exp(z_{\texttt{NO}})}$ (2)
@@
Here, $z_{\texttt{YES}}$ denotes the logit corresponding to the token YES. The probability $P(\texttt{YES})$ then serves as the model’s confidence in semantic equivalence for the given concept pair. Concept pairs with $P(\texttt{YES})$ exceeding a predefined threshold are retained for alignment.
Using token-level likelihoods as confidence estimates has been shown to correlate with model correctness in controlled-output settings, particularly when the answer space is restricted to a small set of discrete labels [35, 36]. In our case, the mapping judgement is explicitly formulated as a binary decision (YES/NO), making token probability a natural and computationally efficient confidence proxy. This probability-based scoring approach substantially reduces inference time, eliminates the need for generating full-length textual responses, and provides a straightforward decision rule while maintaining high alignment precision.
@@
Table 4: The prompt template used for binary equivalence classification between ontology concepts. The fields in the brackets in the user message are populated with structured ontology information.
5)Post-processing and Result Fusion
To ensure the quality and reliability of the final alignment output, we apply a post-processing stage to filter and refine the candidate correspondences generated by the LLM. This stage relies on two thresholds, $\lambda_{\text{prob}}$ and $\lambda_{\text{cs}}$, which control confidence-based and similarity-based filtering, respectively.
Specifically, a threshold $\lambda_{\text{prob}}$ is imposed on the token-level probability associated with the YES response produced by the LLM. Candidate pairs whose confidence scores fall below this threshold are discarded. In parallel, cosine similarity scores obtained during candidate generation are considered, and pairs with embedding similarity lower than $\lambda_{\text{cs}}$ are removed to prevent semantically distant matches from being retained.
After this dual-filtering step, to further enhance precision, the remaining LLM outputs are merged with the results of two exact matching systems, LogMapLt [^4] https://github.com/ernestojimenezruiz/logmap-matcher and BERTMapLt [^5] https://github.com/KRR-Oxford/DeepOnto/tree/main/src/deeponto/align/bertmap. These systems are lightweight variants of their original models and are used here solely for string-based matching.
The detailed configuration of $\lambda_{\text{prob}}$ and $\lambda_{\text{cs}}$ are further analysed in Sections 4 and 6.
4 Overall System Performance Evaluation